A ball is kicked at an angle of 35 with the ground . What is the initial velocity of the ball so that it hits a target that is 30 metres away at a height of 1.8 metres?


1
Expert's answer
2021-09-06T11:33:16-0400

Given:

"\theta=35^{\circ}"

"x=30\:\rm m"

"y=1.8\:\rm m"


The trajectory equation of a ball

"y=x\tan\theta-\frac{gx^2}{2v_0^2\cos^2\theta}"

We get

"1.8=21-6572\/v_0^2"

Finally,

"v_0=\sqrt{342}=19\:\rm m/s"